OpenStudy (anonymous):

differentiate in respect to x x^2/x-1 @mark_o.

OpenStudy (anonymous):

v Du - u Dv use the formula D(u/v)= ---------------- v^2

OpenStudy (anonymous):

yeah i did lol but i think i stuffed something up, can you please solve it for me

OpenStudy (anonymous):

let u=x^2, u= x-1, then (x-1)2x - x^2 ----------------- (x-1)^2 can you continue that pls...

OpenStudy (anonymous):

im sorry v=x-1, u=x^2

OpenStudy (anonymous):

@sirm3d, do you have something to help yasmin plzz? you can go ahead and reply ... :D

OpenStudy (sirm3d):

\[\large v=x-1\\\large Dv = 1 - 0 = 1\\\large u=x^2\\\large Du=2x\] just sub the expressions u, Du, v, and Dv in \[\large D\left(\frac{u}{v}\right)=\frac{v Du - u Dv}{v^2}\] as given by @mark_o.